method of substitution to solve the system of linear equation
6a=7-b
6a=3-b

Answer by TimothyLamb(4379)

L1: 6a = 7 - b
L2: 6a = 3 - b
---
L2: 6a = 3 - b
b = (3 - 6a)
---
L1: 6a = 7 - b
6a = 7 - (3 - 6a)
6a = 7 - 3 + 6a
0 = 4
---
Answer:
The lines L1 and L2 do not intersect.
---
Check the answer by converting L1 and L2 to slope-intercept form: y = mx + b
---
L1: 6a = 7 - b
a = 7/6 + (-1/6)b
a = (-1/6)b + 7/6
---
L2: 6a = 3 - b
a = 3/6 + (-1/6)b
a = (-1/6)b + 3/6
---
Now we can see that L1 and L2 have the same slope (-1/6).
But L1 and L2 have different y-intercepts (7/6 for L1 and 3/6 for L2).
So L1 and L2 are parallel lines that never intersect.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php